# Two axes in non-associative algebras with a Frobenius form

### v4

@@ Abstract.
Throughout this paper is a commutative non-associative algebra over a field of characteristic not In addition posses a Frobenius form. We obtain detailed information about the multiplication in given two axes of type half in
@@
Primary: 17C27, 17A99, Secondary: 17C101. Introduction
Throughout this paper is a commutative non-associative algebra over a field of characteristic not Assume that is a Frobenius form on that is, is a symmetric bilinear map such that We denote by the radical of . Recall that for we denote We direct the reader to Definition 1.1 below and the beginning of Β§2 for definitions required for the introduction.
@@
Throughout the paper, except for Β§2, when we write (but there is no element around) we just mean
@@
The following result is essentially a consequence of [D, Theorem 1.2 and Proposition 4.3].
@@
Theorem 1.5 (Theorem 5.3).
@@
for all
For cases , we have the following theorems.
Theorem 1.7 (Theorem 7.1).
@@
Theorem 1.8.
We were not able to decide what happens when and (though we know that ). There is a good number of additional identities obtained in this paper. We left the case for future works.
@@ 2. The polynomials and
Hence if then β
Lemma 2.2 (Serssβ Lemma).
@@
(b):Β Assume that Let and By Lemma 2.1, Also since is Frobenius, Hence
@@ 3. Derivations
A derivation on is a -linear map that satisfies Leibnizβs law:
@@
(1)Β First note that hence
@@
(4) This is immediate from (1), (2) and (3). β
@@ 4. Obtaining identities in
Let be an axis of Jordan type half. Then
and
and
Remark 4.5.
Lemma 4.6.
Putting in Lemma 4.4 gives
Proof.
(3) Using (2) and Lemma 4.1(4),
Theorem 4.7.
Let and then
Proof.
Lemma 4.8.
Putting and in Lemma 4.4 we get
Hence
Proof.
Hence
This shows (1) and similarly (4) holds.
Hence, by (2),
Multiplying by using and we get
Now we use to see that
so after cancellation we are left with β
Lemma 4.9.
Putting and in Lemma 4.4 we get
Proof.
By Lemma 4.4, replacing with and with
Hence, since the -component is
And, since the -component plus the -compoments are
Lemma 4.10.
For all
Proof.
5. Main theorems
In this section we prove our main results. As before, are axes.
Theorem 5.1.
If then
Proof.
We first show
Theorem 5.2.
Assume then
For all
For all
If then for all
If then either or
Proof.
(1):Β Using Lemma 4.8(4) and Lemma 4.2(10b) we get
(using Theorem 5.1) | ||
(using Theorem 4.7(3)) | ||
(2): Follows from (1) and Lemma 4.6(3), since
(3): Let and Recall from Theorem 4.7(6) that and similarly for By (1) we get,
We conclude that
Multiplying (5.2) by and using (2) we see that
This is because by (2),
Assume Notice that since for all we see that with for Hence, by (5.3),
Assume furthermore that Taking (so that ) we see that since we conclude that
As a corollary, we get
Theorem 5.3.
Assume then
for all
Proof.
(1):Β By Theorem 4.7(4a), by Theorem 5.2 (3), Therefore, for some Multiplying by and recalling that we see that so and (1) hold.
(2): By Theorem 4.7 (8), Therefore, by (1),
(3):Β By Theorem 4.7(6), so, by (1),
(4):Β By Theorem 4.7(3), By (3), Therefore, β
Theorem 5.4.
If then
Proof.
Theorem 5.5.
If then
As an immediate corollary we have
Theorem 5.6.
If then is a derivation.
6. The case
In this section we assume that are axes and that We denote by the radical of the Frobenius form.
Theorem 6.1.
Proof.
First note that by [HSS, Lemma 3.1.9], By Lemma 4.2(7), Theorem 5.1, and Theorem 4.7(7), for all
Multiplying (6.1) by using Theorem 4.7 (1) and (8) we get
Replacing with in (6.1) using Theorem 4.7 (2) and (7) and (6.2), we get
Multiplying (6.3) by using (6.2) we now get
so the theorem holds. β
Theorem 6.2.
If then is a derivation.
Assume Then
Assume Then
In particular, is a derivation.
Proof.
Note also that
(2) (A) We first show that
Let By Lemma 4.10(1),
We now use Theorem 4.7 and By Theorem 4.7(1), and by Theorem 4.7(5), Thus, changing with in (6.6) we get,
By Lemma 4.10(2),
Putting in place of and using again that and (6.5), we get
So multiplying by yields Since (A) holds.
Lemma 6.3.
(6.9) -
Proof.
(1) Replacing in (6.6) by we get
Lemma 6.4.
We can now prove part (2B) of the theorem. By Lemma 4.9(2)
So part (Bi) follows from Lemma 6.4(4ii), since
(C) We start this part by showing
Lemma 6.5.
Proof.
Putting in place of in (6.7), we get
Now part (2C) follows from (6.7), since and as
7. The case
Recall that denotes the radical of the Frobenius form. In this section we prove
Theorem 7.1.
If then is a derivation.
We prove the theorem by showing that assertions (a), (b) and (c) of Theorem 3.7(ii) hold. Assertion (a) holds by Theorem 5.1. Theorems 7.2, 7.3 and 7.4 deal with assertions (b) and (c).
7.1. The case where
In this subsection we assume that Note that is an ideal. Note also, that if then because In particular, if then
Theorem 7.2.
Assume then
Proof.
By Lemma 4.10(2)
Replacing with in (7.1) using Theorem 4.7(5), we get
because by Theorem 4.7(1),
By Lemma 4.10(1),
Theorem 7.3.
Let and then
Proof.
Using also Theorem 4.7, we have
By Lemma 4.9(2)
and by Lemma 4.3(3)
We prove the theorem with a series of short claims.
CLAIM 1.
Proof.
CLAIM 2.
CLAIM 3.
Proof.
(1):Β Replacing in CLAIM 2, with we get
Hence
Multiplying by gives (1).
(2): Replacing in (7.10) with using and CLAIM 1(2),
CLAIM 4.
Proof.
(1):Β By CLAIM 3(1), and by Multiplying CLAIM 1(1) by we get
CLAIM 5.
Proof.
This follows from Lemma 4.9(1), Also using CLAIM 4(2). β
CLAIM 6.
Proof.
Multiplying (7.11) by using CLAIM 3(2),
CLAIM 7.
Proof.
This follows from CLAIM 2 and CLAIM 6. β
CLAIM 8.
Proof.
Part (1) follows from (7.10) and CLAIM 1(1), and (2) follows from (1) and CLAIM 6. β
Notice now that the theorem is immediate from CLAIM 7 and CLAIM 8(2). β
7.2. The case where
Theorem 7.4.
Let and then
Appendix A
In this appendix, we put two lemmas that are not being used in this paper but may find applications in the future.
Lemma A.1.
@@ 7.2. The case where
Lemma A.2.

### v5

@@ Abstract.
Throughout this paper is a commutative non-associative algebra over a field of characteristic not In addition possesses a Frobenius form. We obtain detailed information about the multiplication in given two axes of type half in
@@
Primary: 17C27, 17A99, Secondary: 17C10Contents
1. Introduction
Throughout this paper is a commutative non-associative algebra over a field of characteristic not Assume that is a Frobenius form on that is, is a symmetric bilinear map such that We denote by the radical of . Recall that for we denote We direct the reader to the definition 1.1 below and the beginning of §2 for the definitions required for the introduction.
@@ 1. Introduction
Indeed, Theorem 1.4 introduces new identities, and there are many additional identities in this paper which we believe will become very useful, in particular, for giving an elementary proof that finitely generated three transposition groups are finite. Theorem 1.2 is new as well.
@@
Throughout the paper, except for §2, when we write (but there is no element around) we just mean
@@
The following result is proven in §5, and is essentially a consequence of [D, Theorem 1.2 and Proposition 4.3], but here we prove it using elementary means.
@@
Theorem 1.5 (which infact are two separtate theorems) is proven in §6 and §7 by elemntary means, compare with [D, Theorem 1.2 and Proposition 4.3]).
Theorem 1.5 (Theorems 6.1 and 7.1 respectively).
Suppose then
In particular, is a derivation.
For case , we have the following theorem, proven in §8 by elemntary means.
Theorem 1.6.
@@
If then idetities (1) and (2) of Theorem 1.5 hold. In particular, is a derivation.
We left the case for future works.
@@ 2. The polynomials and
Hence if then ∎
Lemma 2.2 (Serss’ Lemma).
@@
(b): Assume that Let and By Lemma 2.1, Also since is Frobenius, Hence
@@
The following lemma will be used in the next section.
@@ 3. Derivations
A derivation on is a -linear map that satisfies Leibniz’s law:
@@
(1) First note that hence
@@
(4) This is immediate from (1), (2) and (3). ∎
@@ 4. Obtaining identities in
((3)) Taking and in (2) we get,
This shows (3).
(4) (a): Replacing and in (2) we get
Hence, comparing the -components, we get
Hence
or
@@
(5) Taking and in (2) we get
Writing and we get so
Now so part (a) holds. Then, (b) follows from (a) and Lemma 4.1(5)(a).
@@
(8) Interchanging and in (7) we get, by subtracting,
The other part of (8) is clear.
(9) Comparing the -components in (6) we get,
(10) This is immediate from (9). ∎
@@
(2) Comparing the -components in (1) we get
This shows (2).
(3) Comparing the -component plus the -components on both sides of (1) gives:
@@
(2) Comparing the -components we get,
(3) This follows from comparing the components of both sides of (2). ∎
Lemma 4.5.
@@
If then
If then
Proof.
(2) Comparing the -component in (1) we get
(3) This is obvious.
(4) Comparing the -components in (1) we get
(5) Interchanging in (4) we get,
(6) This is obvious.
Lemma 4.6.
Let be an axis of Jordan type half. Then
and
and
Remark 4.7.
Lemma 4.8.
Putting in Lemma 4.6 gives
Theorem 4.9.
Let and then
Proof.
Lemma 4.10.
, Putting and in Lemma 4.6 we get
Hence
Proof.
Hence
This shows (1) and similarly (4) holds.
Hence, by (2),
Multiplying by using and we get
Now we use to see that
so after cancellation we are left with ∎
Lemma 4.11.
Putting and in Lemma 4.6 we get
Proof.
By Lemma 4.6, replacing with and with
Hence, since the -component is
And, since the -component plus the -compoments are
Lemma 4.12.
Proof.
(1) By Theorem 4.9(1),
(2) Replacing with in (1), and using Theorem 4.9(1), we see that
(3) is clear and (4) follows from Theorem 4.9(1). (5) follows from (3) and (4) replacing with Similarly (6) follows from (4). For (7) we have and (8) follow from (7).
(9)&(10) This shows (9), and (10) is obtained by replacing in (9) with ∎
Lemma 4.13.
For all
(4.1) -
Hence,
If then
If then,
5. The proof that
In this section we prove.
Theorem 5.1.
If then
Proof.
We first show the following.
6. The proof of the identity
In this section we prove,
Theorem 6.1.
If then
We start with,
Lemma 6.2.
If then
If then
If then
7. The proof of the identity
The purpose of this section is to prove:
Theorem 7.1.
If then
Proposition 7.2.
Proof.
Lemma 7.3.
Proposition 7.4.
Proof.
Lemma 7.5.
If then
Proof.
(1) Putting in (7.5) in place of we get
(3) By Lemma 6.2(1),
so the first equality in (3) follows from by replacing with and with The second equality follows from this and (1). ∎
Corollary 7.6.
If then
Proof.
(2) Follows from (1) and Lemma 7.5(3). ∎
Corollary 7.7.
Proposition 7.8.
Let If then
Proof.
We can now prove Theorem 7.1.
8. The case
In this section we assume that are axes and that
Theorem 8.1.
If then is a derivation.
Assume Then
In particular, is a derivation.
Proof. (1) Suppose Then, by Lemma 4.3(3), By Lemma 4.13(2), because So (1) follows from Theorem 3.7.
(2) Note that since theorem 4.9 implies
Lemma 8.2.
Proof.
so (1) holds.
(2) This follows by interchanging and in (1) and adding.
(3) Multiplying (1) by using (8.1) we get
(6) By (2) and (5),
so (6) holds. ∎
For the proof of part (2ii), we need the following three lemmas.
Lemma 8.3.
Proof.
(3) Replacing in (1) with we get
(4&5) By Lemma 4.3(3),
Replacing with , and using (8.1), we get
so (4) holds; going back to () we get (5).
(7) (i) Multiplying (5) by using (8.1) we get
(ii) This follows from (7.1) and (i).
(iii) This follows from (i).
Lemma 8.4.
Proof.
This follows from equation (7.2). ∎
Lemma 8.5.
If then